OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4. slope = -1; y intercept at (0, -4/3) slope = 1; y intercept at (0, -4/3) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (anonymous):

slope = 1, y- intercept = (0, -4/3)

OpenStudy (anonymous):

3x-4y = 4 3x-4=4y y=x-4/3

OpenStudy (anonymous):

Thanks soo much

OpenStudy (anonymous):

welcome!!